# Optimization in quantum metrology and entanglement theory using semidefinite programming

### v3

@@ V.4 Maximization of the Hilbert-Schmidt norm
where is a density matrix, are operators, and are constants. The constraints require that the expectation values of the operators equal the constants The optimization problem will lead to pure states if there are pure states satisfying the condition. It can be solved with the method of moments as well as with a see-saw iteration. The see-saw method shows a rapid convergence. In fact, it converges typically in a single step. After examining this simple case, we can have more complicated constraints on the quantum state. For instance, we can consider PPT states or states with a given negativity [10].
@@ V.5 Maximization of the trace norm
The CCNR criterion can detect weakly entangled quantum states, called bound entangle states, that have a positive partial transpose. From this point of view, it is a very relevant question to search for PPT states which violate Eq. (79) the most. This can be done by maximizing over PPT states. We carried out a numerical optimization, using semidefinite programming. We present the maximum of for PPT states for various dimensions in Table 2. The state for the case could be determined analytically, see Appendix E.
@@ VI Conclusions
We discussed how to maximize the metrological gain over local Hamiltonians for a bipartite quantum system. We presented methods that exploit the fact that the the quantum Fisher information can be written as a bilinear expression of the elements of the Hamiltonian, if the Hamiltonian is given in the basis of the density matrix eigenstates. We showed an iterative see-saw (ISS) method that can maximize the quantum Fisher information for large systems. We showed another method based on quadratic programming. We tested the two approaches on concrete examples. We also considered several similar problems that can be solved based on such ideas, for instance, obtaining the maximum violation of the Computable Cross Norm-Realignment (CCNR) criterion for states with a positive partial transpose. Our work can straightforwardly be generalized to multiparticle systems.
@@ VII Acknowledgements
We thank I. Apellaniz, R. Augusiak, R. Demkowicz-Dobrzański, O. Gühne, P. Horodecki, R. Horodecki, and J. Siewert for discussions. We thank J. Kołodyński, K. Macieszczak, and R. Demkowicz-Dobrzański for discussing the details of Ref. [9]. We acknowledge the support of the EU (COST Action CA23115, QuantERA MENTA, QuantERA QuSiED), the Spanish MCIU (Grant No. PCI2022-132947), the Basque Government (Grant No. IT1470-22), and the National Research, Development and Innovation Office NKFIH (Grant No. 2019-2.1.7-ERA-NET-2021-00036). We acknowledge the support of the Grant No. PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by "ERDF A way of making Europe". We thank the "Frontline" Research Excellence Programme of the NKFIH (Grant No. KKP133827). We thank the Project No. TKP2021-NVA-04, which has been implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme. We thank the Quantum Information National Laboratory of Hungary. G. T. thanks a Bessel Research Award of the Humboldt Foundation. We also acknowledge the support of the EU (QuantERA eDICT, CHIST-ERA MoDIC) and the National Research, Development and Innovation Office NKFIH (No. 2019-2.1.7-ERA-NET-2020-00003, No. 2023-1.2.1-ERA_NET-2023-00009, No. KH125096, and No. K145927).
@@ Appendix A Special set of separable states similar to the set of local Hamiltonians
However, if is not known, the recovery is not unique. Eqs. (86) yield a solution for any , for which and holds, constraining as
@@
holds for all Thus, the number of nonzero elements is at most
@@ Appendix C Alternative method to optimize quadratic functions
on a convex set. We could check the extreme point of the set, however, we look for a method that can easily be generalized for larger number of variables.
@@ Appendix D Optimization over the state rather than the Hamiltonian
Then, the symmetric logarithmic derivative, defined in Eq. (15), multiplied by is of the form
@@
In the more general case of larger systems, the quantum Fisher information is not the quadratic function of the density matrix elements anymore [111, 112]. However, even in this case, if is maximizing the quantum Fisher information then
@@
holds, where is the maximal eigenvalue of given in Eq. (132). Thus, every steady state must fulfill the equation of the given Eq. (135) for some Let us consider first the case that has a non-generate spectrum. Let us look for the eigenvectors of the operator
@@
Thus, the mechanism of the see-saw resembles the power iteration discussed in Sec. V.3, however, now we maximize the variance of an operator. It has been shown that the iteration leads to states with a variance that does not decrease [9]. Together with our argument that the steady states all maximize the variance, we can see that the iteration always finds the global maximum. We verified this claim based on extensive numerical tests 222In principle, it is possible to find an initial state for which calculated by Eq. (15), equals the zero matrix. This is the case if is not in the range of the initial density matrix. However small random added noise can easily help, and the iteration will lead to the global maximum..
@@
c. f. Eq. (135). After a derivation similar to the previous one, we find again that maximizing and the same states maximize it as before. Thus, based on the same arguments, the algorithm always finds the global maximum.
@@
The inequalities in Eq. (149) can be added as conditions to an optimization using semidefinite programming. Alternatively, we can also define the a set of quantum states via the physical map, e. g.,
@@
More practical, relevant subsets of physical quantum states are PPT states in a bipartite system. Metrology with such states have been studied using semidefinite programming [10]. A family of bipartite PPT entangled states have been found numerically maximizing the quantum Fisher information [10]. Here, semi-definite programming was used to maximize over the set of PPT quantum states. Later, PPT states having the same quantum Fisher information have been found analytically, which is a strong indication that the method found the global maximum [12].
@@
Appendix E Bound entangled state violating the CCNR criterion maximally

### v4

@@ V.4 Maximization of the Hilbert-Schmidt norm
where is a density matrix, are operators, and are constants. The constraints require that the expectation values of the operators equal the constants The optimization problem will lead to pure states if there are pure states satisfying the condition. The optimization can be solved with the see-saw method as follows
where and are density matrices. The see-saw method shows a rapid convergence. In fact, it converges typically in a single step. The problem can also be solved with the method of moments. After examining this simple case, in addition to the linear constraints, we can have more complicated constraints on the quantum state. For instance, we can consider PPT states or states with a given negativity [10].
@@ V.5 Maximization of the trace norm
The CCNR criterion can detect weakly entangled quantum states, called bound entangled states, that have a positive partial transpose. From this point of view, it is a very relevant question to search for PPT states which violate Eq. (87) the most. This can be done by maximizing over PPT states. We carried out a numerical optimization using semidefinite programming (e. g., optimization over PPT states has been considered in Refs. [80, 10]). We present the maximum of for PPT states for various dimensions in Table 2. The state for the case could be determined analytically, see Appendix F.
@@ VI Conclusions
We discussed how to maximize the metrological gain over local Hamiltonians for a bipartite quantum system. We presented methods that exploit the fact that the quantum Fisher information can be written as a bilinear expression of the elements of the Hamiltonian, if the Hamiltonian is given in the basis of the density matrix eigenstates. We showed an iterative see-saw (ISS) method that can maximize the quantum Fisher information for large systems. We showed another method based on quadratic programming. With this so-called relaxation method we can confirm whether the maximum found by the ISS method is indeed a global one for small systems.
We tested the two approaches on concrete examples, which include one and two copies of the isotropic state of two -dimensional qudits and some important bound entangled states, like, for instance, the Horodecki state. These examples also show that we have provided a simple, systematic and efficient method for maximizing the metrological performance (i.e., the metrological gain) of bipartite quantum states with local Hamiltonians. Our work can straightforwardly be generalized to multiparticle systems.
We also considered several similar problems that can be solved based on such ideas, for instance, obtaining the maximum violation of the Computable Cross Norm-Realignment (CCNR) criterion for states with a positive partial transpose, maximizing various norms, searching for the extremal eigenvalues of Hermitian matrices, and also maximizing the Wigner-Yanase skew information. Thus, we demonstrated that ISS methods can be used beyond metrological performance optimization in other important quantum information applications.
@@ VII Acknowledgements
We thank I. Apellaniz, R. Augusiak, R. Demkowicz-Dobrzański, O. Gühne, P. Horodecki, R. Horodecki, and J. Siewert for discussions. We thank J. Kołodyński, K. Macieszczak, and R. Demkowicz-Dobrzański for discussing the details of Ref. [12]. We acknowledge the support of the EU (COST Action CA23115, QuantERA MENTA, QuantERA QuSiED), the Spanish MCIU (Grant No. PCI2022-132947), the Basque Government (Grant No. IT1470-22), and the National Research, Development and Innovation Office NKFIH (Grant No. 2019-2.1.7-ERA-NET-2021-00036, Advanced Grant No. 152794). We acknowledge the support of the Grant No. PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by "ERDF A way of making Europe". We thank the "Frontline" Research Excellence Programme of the NKFIH (Grant No. KKP133827). We thank the Project No. TKP2021-NVA-04, which has been implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme. We thank the Quantum Information National Laboratory of Hungary. G. T. thanks a Bessel Research Award of the Humboldt Foundation. We also acknowledge the support of the EU (QuantERA eDICT, CHIST-ERA MoDIC) and the National Research, Development and Innovation Office NKFIH (No. 2019-2.1.7-ERA-NET-2020-00003, No. 2023-1.2.1-ERA_NET-2023-00009, No. KH125096, and No. K145927).
@@ Appendix A Special set of separable states similar to the set of local Hamiltonians
However, if is not known, the recovery is not unique. Eqs. (94) yield a solution for any , for which and hold, constraining as
@@
holds for all Thus, the number of nonzero elements is at most The remaining elements of the coherence vector are given as
The following statement can straightforwardly be proven. If the state fulfills the condition in Eq. (103) then it can be written in the form given in Eq. (88). Thus, we can decide whether a state is of the form given in Eq. (88) based on verifying that Eq. (103) holds.
@@ Appendix C Alternative method to optimize quadratic functions
on a convex set. We could check the extreme point of the set, however, we look for a method that can easily be generalized to a larger number of variables.
@@ Appendix D Optimization over the state rather than the Hamiltonian
Then, we set to the symmetric logarithmic derivative defined in Eq. (15) to
@@
In the more general case of larger systems, the quantum Fisher information is not a quadratic function of the density matrix elements anymore [119, 120]. However, even in this case, if is maximizing the quantum Fisher information then
@@
holds, where is the maximal eigenvalue of given in Eq. (141). Thus, every steady state must fulfill the condition given in Eq. (144) for some Let us consider first the case that has a non-degenerate spectrum. Let us look for the eigenvectors of the operator
@@
Thus, the mechanism of the see-saw resembles the power iteration discussed in Sec. V.3, however, now we maximize the variance of an operator. It has been shown that the iteration leads to states with a variance that does not decrease [12]. Together with our argument that the steady states all maximize the variance, we can see that the iteration always finds the global maximum. We verified this claim based on extensive numerical tests 222In principle, it is possible to find an initial state for which calculated by Eq. (15), equals the zero matrix. This is the case if is not in the range of the initial density matrix. However, small random added noise can easily help, and the iteration will lead to the global maximum..
@@
c. f. Eq. (144). After a derivation similar to the previous one, we find again that the algorithm is maximizing and the same states maximize it as before. Thus, based on the same arguments, the algorithm always finds the global maximum.
@@
The inequalities in Eq. (158) can be added as conditions to an optimization using semidefinite programming. Alternatively, we can also define a set of quantum states via the physical map, e. g.,
@@
More practical, relevant subsets of physical quantum states are PPT states in a bipartite system. Metrology with such states have been studied using semidefinite programming [10]. A family of bipartite PPT entangled states have been found numerically maximizing the quantum Fisher information [10]. Here, semidefinite programming was used to maximize over the set of PPT quantum states. Later, PPT states having the same quantum Fisher information have been found analytically, which is a strong indication that the method found the global maximum [11].
@@
Appendix E An example for using the moment method to upper bound the quantum Fisher information
In this Appendix, we exemplify the use of the moment method discussed in Sec. IV.2 to upper bound the quantum Fisher information for a fixed state . We will focus on , however, the method generalizes straightforwardly to any .
Let us first recap the optimization problem. Fix a state with the eigendecomposition as in Eq. (9). According to Eq. (8), our task is to maximize
over bipartite local Hamiltonians of the form given in Eq. (24), where we impose the condition
The matrix elements of in the eigenbasis of are given by . Let us denote by the maximum value in Eq. (162) over local Hamiltonians. Note that the above condition is a relaxation of the stricter one given in Eq. (64) for with . Therefore, in general defines only an upper bound to the true maximum of the quantum Fisher information. The constant coefficients depend only on the eigenvalues of the probe state and are given by Eq. (10) whenever the denominator is nonzero.
Now we apply the following steps to bring the problem to an SDP.
(i) Build an orthonormal basis of Hermitian matrices. For , we choose the four normalized Pauli matrices , where is the identity, and the rest are , , and .
(ii) Represent the local Hamiltonian operators as
with real coefficients , , forming the vector . With this choice we have
that is, we recover the condition in Eq. (163).
(iii) For each , define the matrices and , and from them construct the complex vector compactly as
(iv) In terms of we can write the complex matrix element as
(v) Hence the quantum Fisher information is
where the operator
is a real symmetric matrix, and we used that is real valued.
This leads us to the optimization problem
| subject to |
Thus, we arrived at an optimization task of the form given in Eq. (54). However, we can reformulate this problem as an SDP
| subject to | ||||
where we can identify .
Dropping the non-convex rank constraint yields a convex SDP [see also Eqs. (62) and (63)]. However, it can be shown that the optimum of this relaxed SDP is attained at the rank-one matrix , where is any unit eigenvector of associated with the largest eigenvalue [97, 98]. Hence the solution to the task (170) is , and the optimal objective value is
As an illustrative example, let us apply the above procedure to calculate for the two-qubit isotropic state from Eq. (61) (see also the first line of Table 1). In this case, the matrix from Eq. (169) evaluates to
The eigenvalues of this symmetric matrix are
Hence, by Eq. (172), we obtain
One of the eigenvectors corresponding to the maximal eigenvalue can be written as
where and . Note that since is three-fold degenerate, any unit vector in the corresponding eigenspace gives an optimal .
From this choice we obtain , and by Eq. (164) the optimal Hamiltonian is given by the bipartite Hamiltonian given in Eq. (24) with
Incidentally, we find , which also satisfies the original (non-relaxed) constraint of Eq. (64). Thus, in this case the upper bound in (175) is achievable, and the bound is tight.
Appendix F Bound entangled state violating the CCNR criterion maximally